Find the minimum and maximum values of z=9x+22​y, if​ possible, for the following set of constraints.
4x+3y≥12
x+3y≥6
x≥​0, y≥0

Answer by greenestamps(13200)



I don't know if this is supposed to be a duplicate submission of a problem, or if is a corrected version; the other one had an objective function of 9x+2y; this one shows an objective function of 9x+22y.

Now that I look at the two problems more closely, I see that it would make sense to have the same constraints but different objective functions in the two problems. So they probably were supposed to be two different problems.

With the new objective function, the response will be very nearly the same... except that the new objective function gives a minimum value at a different vertex.

Graph the equations of the boundary lines of the inequalities.

4x+3y=12 has intercepts (3,0) and (0,4).
x+3y=6 has intercepts (6,0) and (0,2).
The intersection of the two lines is at (2,4/3)

Since both inequalities have "y >= ...", the feasibility region is unbounded; the three vertices of the feasibility region are (0,4), (2,4/3), and (6,0).

Evaluate the objective function at each vertex of the feasibility region.
Note that there will be a minimum value of the objective function; but with the unbounded feasibility region, there will be no maximum value.

(0,4): 9(0)+22(4) = 0+88 = 88
(2,4/3): 9(2)+22(4/3) = 18+88/3 = 142/3 = 47 1/3
(6,0): 9(6)+22(0) = 54+0 = 54

Now the minimum value of the objective function subject to the given constraints is 47 1/3, at (2,4/3).

Note that it is not necessary to evaluate the objective function at every vertex of the feasibility region. In this fairly simple example, it was easy to do so. But in more complicated similar problems, you can get to the answer with much less work without evaluating the objective function at every vertex, as described below.

The objective function is z = 9x+22y. The slope of the graph of every function of this form is -9/22. By looking at even a rough sketch of the feasibility region, you can see that the only place where a line with slope -9/22 just touches the feasibility region, without passing through it, is at the vertex (2,4/3).

Furthermore, you can see that, since the feasibility region is unbounded, there is no place other than at (2,4/3) where a line with slope -9/22 will touch the feasibility region in only one place; that indicates that there is no maximum value to the objective function.

Finally, assuming these two questions are from the same person, note that if the objective function were 9x+50y, then the slope of the objective function line would be -9/50; and then the vertex where the minimum value of the objective function is obtained would be (6,0).